A manufacturer has a policy that a 16 ounce bottle of salad can be off as much as 0.06 ounces off of the advertised weight. Which inequality describes the acceptable volume of a bottle of salad dressing?
A. |x- 16| less than or equal to 0.06
B. x- 16 less than or equal to 0.06

Answer by Theo(13342)

you want selection A.

|x-16| <= .06

this expression can be translated as follows:

(x-16) <= .06 when the expression is positive or zero.

-(x-16) <= .06 when the expression is negative.

multiply both sides of the second expression by -1 and you will get (x-16) >= -.06 because multiplying both sides of an inequality by a negative number reverses the inequality.

so you have (x-16) <= .06 and (x-16) <= .06

if you solve for x in both of these equations, you will get x <= 16.06 and x >= 15.94

this says that the weight of the bottle can be up to .06 ounces more than 16 ounces or up to .06 ounces less than 16 ounces.

the second equation of (x-16) <= .06 only covers the case where the weight of the bottle can be up to .06 ounces more than 16 ounces. it does not cover the case where the weight of the bottle can be up to .06 ounces less than 16 ounces.

the absolute value equation covers both cases.